# SPECIAL RELATIVITY

### v9

@@ 7.5 Aberration of Relativistic Particles
Suppose we have a particle source. All particles which come out of the source will have the same velocity $V$. Particles trace paths along the $z$-axis. We consider the case when the source is at rest in the initial inertial frame and the observer (with his measuring devices) is accelerated from rest up to velocity $v$ along the $x$ axis. It is assumed that $v\ll V$. Suppose that an observer in the accelerated frame performs an aberration measurement. How shall we describe the aberration of particle beam from the source which is at rest in the inertial frame? In order to predict the result of the aberration measurement the accelerated observer should use the Langevin metric Eq.(15). The aberration increment is connected with the problem parameters by the relation $\theta_{a}=-[1+(1-1/\gamma_{p})]\theta_{c}$, where $\theta_{c}=v/V$ is the particle deflection angle according to classical approach. Here $\gamma_{p}=1/\sqrt{1-V^{2}/c^{2}}$ is the relativistic factor. The term $[1+(1-1/\gamma_{p})]$ is new from relativistic (Langevin) metric, but reduced to familiar classical limit as $V^{2}/c^{2}\to 0$. For light $V/c=1$ and there is an extra factor 2. A correct solution of this problem in the rotating frame requires the use of metric tensor even in first-order experiments since the crossed term in the Langevin metric (which involves the first-order deviation of metric tensor from its Minkowski form) plays a fundamental role in the non-inertial kinematics of a relativistic particle propagation. In the Chapter 8 (see Section 8.2) we will find that it is very easy to treat the problem by the use of dynamical line of arguments. The electromagnetic forces which govern the properties of an emitted electron beam must be affected by source acceleration in such a way that they lead to a deviation of the electron transport direction. We can obtain the electromagnetic fields in accelerated frame using Galilean transformation.
We want now to find the aberration of ultrarelativistic particles from stationary and moving sources. Imaging that there are two identical sources. For the sake of simplicity we assume that the relativistic factor of the beam $\gamma_{p}\gg 1$. Let us consider the case when the first source is at rest in an inertial frame and the second source is accelerated from rest up to velocity $v$ along the $x$-axis. Suppose that an observer, which is at rest with respect to the inertial frame of reference performs the direction of the particle transport measurement. Now we must be careful about the initial direction of the velocity of a particle beam. As example, we consider the case in which particles are initially directed along the $z$-axis. Then how does the particle beam from the moving source looks? The inertial observer would find that angular displacement is equal to $\theta_{a}=v/c$. This is example what called aberration of relativistic particles and it is well known.
Figure 40: The aberration of ultrarelativistic ($\gamma_{p}\gg 1)$ particles from stationary and moving sources. If the particle source is at rest relative to the fixed stars and the observer with measuring devices started from rest to uniform motion with velocity $v\ll c$, then the apparent angular position of the source seen in the accelerated frame would jump by angle $\theta_{a}=-2v/c$. This situation is not symmetrical with respect to the change of the reference frames. If observer is at rest relative to the fixed stars and the source started from rest to uniform motion relative to the fixed stars, then the apparent angular position of the source seen in the inertial frame would jump by angle $\theta_{a}=v/c$
Figure 41: Reciprocity in the theory of the aberration of particles. One must use the same beam directing to demonstrate the reciprocity. After the redirection procedure of the accelerated source the inertial observer would find that the angular displacement is $\theta_{a}=2v/c$ (for the sake of simplicity it is assumed that $\gamma_{p}\gg 1$).
The difference between measurements made in both reference frames is very interesting. According to the asymmetry between the inertial and accelerated frames, there is a remarkable prediction on the theory of the aberration of particles. Namely, if the source is at rest relative to the fixed stars and the observer with measuring devices started from rest to uniform motion relative to the fixed stars, then the apparent angular position of the emitter seen in the accelerated frame would jump by angle $\theta_{a}=-2v/c$. This situation is not symmetrical with respect to the change of the reference frames, Fig. 40. It could be said that the crossed term in metric Eq.(15) generates anisotropy in the accelerated frame that is responsible for the difference between aberration increments.
When the accelerated system starts moving with constant velocity the procedure of beam redirection can be performed. Let us describe what happens when accelerated observer performed the redirection of the accelerated source. We consider the case in which finally the particles trace paths along the $z_{n}$-axis. After this redirection procedure the inertial observer would find that angular displacement is $\theta_{a}=2v/c$. Suppose that the accelerated observer also performs an aberration measurement. Fig. 41 shows that the aberration increment is connected with the problem parameters by the relation $\theta_{a}=-2v/c$.
@@
Above in this chapter we demonstrated that aberration of light effect in the accelerated frame, as viewed from the inertial lab frame, presents a well known kinematic effect of special theory of relativity. What about particles? Can we not look at the aberration of particles in the accelerated frame in the same way? As we shall see, the aberration of particles effect in non-inertial frames of reference can be described in an inertial frame within standard special relativity taking advantage of the Wigner rotation theory.
@@
The Wigner rotation is a relativistic kinematic effect, which consists in that a coordinate axes of a non-inertial reference frame is rotating about the axes of a (Lorentz) initial inertial frame. What is the formula for Wigner rotation? It turns out that this is complicated, and it takes a great deal of study and sophistication to appreciate it. (We shall do this in the Chapter 17). We only write it down to give the impression. The expression for the rotation angle in the inertial frame can be presented in vector form $\vec{\theta_{w}}=(1-1/\gamma_{p})\vec{V}\times\vec{v}/V^{2}$, where $\vec{V}$ is the vector of particle velocity in the initial Lorentz frame before acceleration, $\vec{v}$ is the vector of small velocity change ($v\ll V$) due to acceleration, $\theta_{w}$ is the Wigner rotation angle of the spatial coordinate axes of the system comoving with a particle relative to Lorentz lab frame. We shall discuss this expression further in the Chapter 17.
@@
To find the rotation magnitude in the stated problem, we introduce a composition of boosts. Let $S$ be the initial inertial frame of reference, $S_{n}$ an accelerated frame with velocity $\vec{v}$ relative to the inertial frame, and $S^{\prime}$ is a particle comoving frame which moves (after the redirection of the accelerated source) relative to the $S_{n}$ with velocity $\vec{V}$ along the $z_{n}$-axis. Two sequential boosts from the inertial frame $S$ to $S_{n}$ and then to $S^{\prime}$ are equivalent to the boost from $S$ to $S^{\prime}$ and the subsequent rotation. The comoving frame $S^{\prime}$ is rotated through the angle $\theta_{w}$ with respect to the inertial frame $S$. Now we must be careful about direction of rotation. There is a good mnemonic rule. The rule says that the direction of the Wigner rotation in the Lorentz lab frame is the same as the direction of the velocity rotation in the Lorentz lab frame. We can easily understand that the coordinate axes of the particle comoving frame $(x^{\prime},y^{\prime},z^{\prime})$ will be parallel to the coordinate axes of the accelerated frame $(x_{n},y_{n},z_{n})$. As viewed from the inertial frame, the coordinate axes of the accelerated frame $(x_{n},y_{n},z_{n})$ is also rotated through the angle $\theta_{w}$ with respect to the coordinate axes of the inertial frame $(x,y,z)$, Fig. 41. Our results show that we cannot remain with the framework of parallel axes $z_{n},z$ when considering the inertial frame view of particle beam observations of the accelerated (up to the velocity $v$) observer in perpendicular geometry. For light $V/c\to 1$ and the axes of accelerated frame are then rotated with respect to the inertial frame axes by angle $\theta_{w}=v/c$.

### v10

@@ 9 Aberration of Relativistic Particles
We want now to discuss the phenomenon of the aberration of relativistic particles. Suppose we have a particle source at rest in the initial inertial frame. All particles which come out of the source will have the same velocity $V$. Particles trace paths along the $z$-axis. Let us consider the case when the observer (with his measuring devices) and the particle source are accelerated from rest up to velocity $v$ along the $x$ axis. It is assumed that $v\ll V$. Suppose that an observer in the accelerated frame performs an aberration measurement. How shall we describe the aberration of particle beam from the source which is at rest in the accelerated frame? To predict the result of the aberration measurement the accelerated observer should use the Langevin metric Eq.(15) (see Section 9.2). The aberration increment is connected with the problem parameters by the relation $\theta_{a}=-(1-\sqrt{1-V^{2}/c^{2}})v/V$. We should point out that this result is valid for an arbitrary Lorentz factor $\gamma=1/\sqrt{1-V^{2}/c^{2}}$. The term $(1-1/\gamma)v/V$ is new from the relativistic (Langevin) metric but reduced to the familiar classical limit as $V^{2}/c^{2}\to 0$. A correct solution to this problem in the accelerated frame requires the use of metric tensor even in first-order experiments since the crossed term in the Langevin metric (which involves the first-order deviation of metric tensor from its Minkowski form) plays a fundamental role in the non-inertial kinematics of relativistic particle propagation.
9.1 Inertial Frame View of Observations of the Noninertial Observer
@@
We can, of course, analyze this situation in various other ways. Above we demonstrated that the aberration of light effect in the accelerated frame, as viewed from the initial inertial frame, presents a well-known kinematic effect of the special theory of relativity. Can we not look at the aberration of particles in the accelerated frame in the same way? As we shall see, the aberration of particles effect in non-inertial frames of reference can be described in an inertial frame within standard special relativity taking advantage of the Wigner rotation theory.
@@
The Wigner rotation is a relativistic kinematic effect, which consists in that the coordinate axes of a non-inertial reference frame are rotating about the axes of a (Lorentz) initial inertial frame. The expression for the rotation angle in the inertial frame can be presented in vector form $\vec{\theta_{w}}=(1-1/\gamma)\vec{V}\times\vec{v}/V^{2}$, where $\vec{V}$ is the vector of particle velocity in the initial Lorentz frame before acceleration, $\vec{v}$ is the vector of small velocity change ($v\ll V$) due to acceleration, $\theta_{w}$ is the Wigner rotation angle of the proper frame axes in the (Lorentz) inertial frame. We want now to see how this formula for the Wigner rotation works in the case when the observer and particle source are accelerated up to velocity $v$ in the inertial frame.
@@
To find the rotation magnitude in the stated problem, we introduce a composition of Lorentz boosts. Let $S$ be the initial inertial frame of reference, $S_{n}$ an accelerated frame with velocity $\vec{v}=\vec{e}_{x}v$ relative to the inertial frame, and $S^{\prime}$ is a frame which moves relative to the $S_{n}$ with velocity $\vec{V}$ along the $z_{n}$-axis. Two sequential boosts from the inertial frame $S$ to $S_{n}$ and then to $S^{\prime}$ are equivalent to the boost from $S$ to $S^{\prime}$ and the subsequent rotation. The frame $S^{\prime}$ is rotated through the angle $\theta_{w}$ with respect to the inertial frame $S$. Now we must be careful about the direction of rotation. There is a good mnemonic rule. The rule says that the direction of the Wigner rotation in the Lorentz lab frame is the same as the direction of the velocity rotation in the Lorentz lab frame. We can easily understand that the coordinate axes of the frame $(x^{\prime},y^{\prime},z^{\prime})$ will be parallel to the coordinate axes of the accelerated frame $(x_{n},y_{n},z_{n})$. As viewed from the inertial frame, the coordinate axes of the accelerated frame $(x_{n},y_{n},z_{n})$ are also rotated through the angle $\theta_{w}$ with respect to the coordinate axes of the inertial frame $(x,y,z)$. Our results show that we cannot remain with the framework of parallel axes $z_{n},z$ when considering the inertial frame view of particle beam observations of the accelerated (up to the velocity $v$) observer in perpendicular geometry. So we have expression for the aberration increment in the accelerated frame: $\theta_{a}=-\theta_{w}=-(1-1/\gamma)v/V$. For light $V/c\to 1$ and the aberration increment becomes $\theta_{a}=-v/c$.
It is very easy to treat the aberration of particle problems by the use of a dynamical line of arguments. The electromagnetic forces that govern the properties of an emitted electron beam must be affected by source acceleration in such a way that they lead to a deviation of the electron transport direction. We can obtain the electromagnetic fields in an accelerated frame using Galilean transformation.
For those who have already received knowledge about special relativity from well-known textbooks, it may be expedient to begin its aberration of particle (theoretical and experimental) development with the microscopic approach. The point is that one can compute any relativistic quantity directly from the underlying theories of matter without involving relativistic kinematics effects at all. The aberration of particles associated with the transformation from the inertial frame to the accelerated frame may be regarded as a result of the action of a certain force. This problem has no relation to the problem of the ether and is already formulated within the framework of special relativity.
Let us consider what happens if the electron gun, an electron source for a relativistic electron beam, in the initial inertial frame is accelerated from rest up to velocity $v$ along the $x$-axis. The electromagnetic forces that govern the properties of an emitted electron beam must be affected by the source acceleration in such a way that they lead to a deviation of the electron transport direction. There is a machinery behind the electron momentum $\vec{p}$ changes when the source and observer (with his measuring instruments) are at rest in the accelerated frame. Its origin is explained in the framework of the electrodynamics theory.
We can obtain the electrodynamics equations in an accelerated reference frame using the Galilean transformation. When we consider the electrodynamics equations under Galilean transformation, there is a question: what is the transformation law for the electromagnetic fields $\vec{E}$ and $\vec{B}$? The fields $\vec{E}_{n}$ and $\vec{B}_{n}$ observed in the accelerated frame where the particle source is at rest are not the same as those $\vec{E}$ and $\vec{B}$ in the inertial frame before the active Galilean boost.
We need to find the relation between $\vec{E}_{n}$ and $\vec{B}_{n}$ and $\vec{E}$ and $\vec{B}$. To do this, we transform coordinates $(t,x,y,z)$ that would be coordinates of an inertial observer $S$ moving with velocity $-v$ with respect to the observer $S_{n}$, using a Galilean transformation: we substitute $x_{n}=x-vt$, while leaving time unchanged $t_{n}=t$ into the Minkowski metric $ds^{2}=c^{2}dt^{2}-dx^{2}-dy^{2}-dz^{2}$ to obtain Langevin metric $ds^{2}=c^{2}(1-v^{2}/c^{2})dt_{n}^{2}-2vdx_{n}dt_{n}-dx_{n}^{2}-dy_{n}^{2}-dz_{n}^{2}$. We can obtain the electrodynamics equations in an accelerated frame using the Galilean transformation (with velocity $-v$) of Maxwell’s equations. With the Galilean transformation, the field transformation, for the case $\vec{B}=0$, is $\vec{E}_{n}=\vec{E}$, $\vec{B}_{n}=-\vec{v}\times\vec{E}/c$. This expression describes first-order (in $v/c$) effects only. Here $\vec{v}$ is the velocity vector of the source in the initial inertial frame.
Figure 38: Aberration of electrons when the electron source and observer are at rest in the accelerated frame.
[…]
We see that the aberration of particles associated with the transformation from the inertial coordinate system to the accelerated coordinate system may be regarded as a result of the action of Lorentz force. The electromagnetic forces that govern the properties of an emitted electron beam must be affected by acceleration with respect to the fixed stars in such a way that they lead to the deviation of the electron beam transport direction, Fig. 38.
The formula $\vec{\theta}_{a}=-(1-1/\gamma)\vec{V}\times\vec{v}/V^{2}$ correspond to the result we demonstrated for observation of the non-inertial observer taking advantage of the Wigner rotation theory. The theory of special relativity shows that the aberration of particles effect in the accelerated frame, as viewed from the initial inertial frame (in the Lorentz coordinatization), presents a kinematic effect of the special theory of relativity. In this case, when viewed from the initial inertial frame, the aberration image shift in the accelerated frame is regulated by the Wigner rotation. According to the Wigner rotation theory, two observers with different trajectories have different 3-spaces. In contrast, the accepted in previous literature incorrect assumption that a non-inertial (e.g. earth-based) observer and an inertial (e.g. sun-based) observer have a common 3-space has always been considered obvious. It is important to stress at this point that the dynamical line of arguments discussed here explains what the Wigner rotation physically means in the rotating frame. The relativistic kinematics effects in the rotating frame, as viewed from the inertial observer, are only an interpretation of the behavior of the electromagnetic fields.
9.3 Magnetic Field Measurements in the Accelerated Electron Source
We would now like to describe an apparent paradox. In special relativity, we are confronted with a few things that might seem counterintuitive. But this is still a mathematically based theory. It tells us a peculiar thing: when we are accelerated electron source, the fields in the accelerated frame are not static; the problem is time dependent. That is what special relativity says! When the electron source and observer are at rest in the accelerated frame the derivative $\partial/\partial t_{n}$ will be nonzero. Intuition would seem to tell us that everything is at rest, so the electric field is not changing with time. But special relativity says that there is a derivative $\partial\vec{E}/\partial t_{n}=v\partial\vec{E}/\partial x_{n}$ because there is $\partial\vec{E}/\partial x_{n}$ that is not zero. This is just a feature of the 4D formalism with no special consequence. A comparison with a gauge transformation in classical electromagnetism might help here. In electrodynamics, it is often convenient to introduce the Coulomb gauge. In this case field propagates instantaneously. But this instantaneous propagation is just a mathematical description in an intermediate step: when a complete calculation is made, proper cancellations of the instantaneous propagation take place.
We obtained the electrodynamics equations in an accelerated frame using the inverse Galilean transformation (with velocity $-v$) of Maxwell’s equations. With the Galilean transformation, the field transformation, for the case $\vec{B}=0$, is $\vec{E}_{n}=\vec{E}$, $\vec{B}_{n}=-\vec{v}\times\vec{E}/c$. The electric $\vec{E}_{n}$ and magnetic $\vec{B}_{n}$ fields are not changing with time - all of the fields are static so that the problem is not time-dependent. There are many physicists who have already received knowledge about electrodynamics from textbooks and who would say, ”There must be currents to get a static magnetic field at all - and currents can only be from moving charges”. Since there are no moving charges , the situation seems indeed paradoxical. Here we have a new kind of situation that is quite different from (initial) inertial frame electrodynamics.
How can that be? When the electron source and observer are at rest in the accelerated frame the derivative $\partial/\partial t_{n}$ will be nonzero. One can easily show this by transforming coordinates $(t,x,y,z)$ that would be coordinates of an inertial observer $S$ moving with velocity $-v$ with respect to the observer $S_{n}$. The variables $x,y,z,t$ can be expressed in terms of the independent variables $x_{n},y_{n},z_{n},t_{n}$ using a Galilean transformation so that fields can be written in terms of $x_{n},y_{n},z_{n},t_{n}$. From the Galilean transformation $x_{n}=x-vt,~{}y_{n}=y,~{}z_{n}=z,~{}t_{n}=t$, after partial differentiation, one obtains $\partial/{\partial t_{n}}=\partial/{\partial t}+v\partial/{\partial x}$, $\partial/{\partial x_{n}}=\partial/{\partial x}$. So $\partial/\partial t_{n}=v\partial/\partial x_{n}$ at $\partial/\partial t=0$.
When the situation is described as we have done here, there doesn’t seem to be any paradox at all; it comes out quite natural that separation between space and time is no longer possible in general. The argument that the result is paradoxical runs something like this: If we could measure magnetic field $\vec{B}_{n}$ by the usual magnetometer there would indeed be a paradox. But we could not do that. We already discussed above that magnetic field $\vec{B}_{n}$ leads to change in the transverse momentum of test particles given by $\theta_{a}=p_{\perp}/p=(1-1/\gamma)v/V$. If a particle motion velocity is non-relativistic, the binomial expansion yields $\vec{\theta}_{a}=-[V^{2}/(2c^{2})]\vec{V}\times\vec{v}/V^{2}$ and the momentum perturbation presents a relativistic effect. So special relativity says that there are no changes in the (test) particle momentum as $V^{2}/c^{2}\to 0$. If we look at the usual (i.e. classical) magnetometer with test particles, we see no evidence of the magnetic field $B_{n}$ effect. Not only does this particular kind of classical magnetometer can not measure the magnetic field $B_{n}$, but if the theory of relativity is correct, any other classical magnetometer, operating on any principle whatsoever, would also appear to demonstrate no evidence of the magnetic field $B_{n}$ effect.
To understand magnetic field measurements in the accelerated electron source, we have to watch the machinery of the classical magnetometer and see what happens when it is at rest in the accelerated frame. Since it is rather difficult, we shall take a simple kind of usual magnetometer. It is relatively easy to observe nuclear magnetism through the phenomena of ”nuclear magnetic resonance”. A proton resonance apparatus can be used as a proton resonance magnetometer. We give a general analysis of this type of classical magnetometer independently of the experimental accuracy. In general, a current loop, which has a magnetic moment $\vec{m}$ in the inertial lab frame in which it is at rest, has an electric dipole moment equal to $\vec{d}=\vec{v}\times\vec{m}/c$ when it is moving with velocity $\vec{v}$ relative to the lab frame. The magnetic properties of materials are attributed to atomic current loops. Then a magnetized body that is accelerated relative to the lab frame should have an electric polarization relative to the lab frame. Thus, in the lab frame, the additional potential energy of a moving particle with the proper magnetic moment $\vec{m}$ in the magnetic field $\vec{B}$ and the electrical field $\vec{E}$ becomes $U_{p}=-\vec{m}\cdot\vec{B}-\vec{d}\cdot\vec{E}$. We obtained the electrodynamics equations in an accelerated frame using the Galilean transformation (with velocity $-\vec{v}$) of Maxwell’s equations. With the Galilean transformation, the magnetic dipole transformation for the case $\vec{d}=0$, is $\vec{m}_{n}=\vec{m}$, $\vec{d}_{n}=-\vec{v}\times\vec{m}/c$. As we know from vector algebra, $(\vec{a}\times\vec{b})\cdot\vec{c}=\vec{a}\cdot(\vec{b}\times\vec{c})$; so our terms are also the same as $(-\vec{v}\times\vec{E})\cdot\vec{m}=-\vec{v}\cdot(\vec{E}\times\vec{m})$, $(-\vec{v}\times\vec{m})\cdot\vec{E}=-\vec{v}\cdot(\vec{m}\times\vec{E})$, as evaluated in the accelerated frame. Hence we obtain the additional potential energy equal to zero just as we expected. We could not measure magnetic field $\vec{B}_{n}$ by usual magnetometer.
9.4 Aberration of Particles from a Moving Source
Let us consider what happens if the electron gun, an electron source for a relativistic electron beam, in the initial inertial frame is accelerated from rest up to velocity $v$ along the $x$-axis, Fig. 39. In other words, we consider a so-called active (physical) boost. In the case of an active boost, we are dealing with the motion of the same physical system, evolving in time and treated from the point of view of the same reference system. The simplest method of synchronization consists in keeping, without changes, the same set of uniformly synchronized clocks used in the case when the particle source was at rest, i.e. we still enforce the clock transport synchronization ( or Einstein synchronization which is defined in terms of light signals emitted by the dipole source at rest). This choice is usually the most convenient one from the viewpoint of the smooth tailoring of the metric tensor. This synchronization convention preserves simultaneity and is based on the absolute time (or absolute simultaneity) convention.
Figure 39: Aberration of electrons from a moving source in an inertial frame. The motion of the electron gun is perpendicular to the optical axis.
We can obtain the electric field $\vec{E}$ and the magnetic $\vec{B}$ observed in the inertial frame where the particle source is moving with velocity $v$ using the Galilean transformation. When we consider the electrodynamics equations under Galilean transformation, there is a question: what is the transformation law for the electromagnetic fields? The equivalence of the active and passive pictures is due to the fact that moving system one way is equivalent to moving the coordinate system the other way by an equal amount. It is for this reason that fields in the inertial frame before the active Galilean boost are the same as those $\vec{E^{\prime}}$ and $\vec{B^{\prime}}$ in the comoving coordinate system. With the Galilean transformation, the field transformation, for the case $\vec{B^{\prime}}=0$, is $\vec{E}=\vec{E^{\prime}}$, $\vec{B}=\vec{v}\times\vec{E^{\prime}}/c$. This expression describes first-order (in $v/c$) effects only. Here $\vec{v}$ is the velocity vector of the source in the lab frame.
[…]
We described the effect of aberration of particles by working only up to the first order $v/c$. An elementary explanation of this effect is well-known. The aberration of particles in the inertial lab frame can be easily explained in terms of the Galilean transformation of velocities between different reference frames. It is important to stress at this point that the dynamical line of arguments discussed here explains what the aberration of particles physically means in the inertial frame. The kinematics effects are only an interpretation of the behavior of the electromagnetic fields.
Now let us return to the magnetic field measurements in the accelerated electron source, Fig. 39. In the lab frame, we are dealing with motion of the electron source, evolving in time and treated from the point of view of the inertial system. It is not difficult to see that the peculiarity of the situation is that here the ”source motion” is a real observable effect. Indeed, the problem is time-dependent. In this situation, the velocity of the source has a physical meaning. In the case of a moving electron source, special relativity says that magnetic field $\vec{B}=\vec{v}\times\vec{E}/c$ leads to a change in the transverse momentum of test particles given by $\theta_{a}=p_{\perp}/p=v/V$ and this result is valid also in the non-relativistic limit $V^{2}/c^{2}\to 0$. Now we can raise an interesting question. Suppose that an inertial observer is measuring the magnetic field with an usual magnetometer. Now we know that there are moving charges, so that means that when there is a current in the moving source, the source generates a magnetic field $\vec{B}=\vec{v}\times\vec{E}/c$ which can be measured with a usual magnetometer. In order to understand this, we have to watch the machinery of the NMR magnetometer. Now that our magnetometer is at rest in the lab frame, that means that the additional potential energy of a proton with the proper magnetic moment $\vec{m}$ in the magnetic field $\vec{B}$ is equal to $-\vec{m}\cdot\vec{B}\neq 0$. The moving electron source in the lab frame generates a magnetic field which can be measured with NMR magnetometer. We come to the conclusion that the magnetic field generated by a moving source in an inertial frame is a real field in the sense we have defined it.